OpenStudy (anonymous):

The length of a rectangle is 5 centimeters less than twice its width. The perimeter of the rectangle is 80 cm. What are the dimensions of the rectangle? length = 23 cm; width = 14 cm length = 22 cm; width = 18 cm length = 25 cm; width = 15 cm length = 24 cm; width = 16 cm

OpenStudy (alexandervonhumboldt2):

look: let a be the lenght and b be the wight. so 2*(a+b)=80 a+b=40 a=2b-5 substitute 2b-5 for a and solve: 2b-5+b=40 3b-5=40 3b=45 b=15 wight a=2*15-5 a=??? lenght

OpenStudy (anonymous):

25?

OpenStudy (alexandervonhumboldt2):

YES so it is C

OpenStudy (anonymous):

Thank youuuuuu :)

OpenStudy (alexandervonhumboldt2):

:)